OpenStudy (anonymous):

A car initially at rest undergoes uniform acceleration for 6.32 seconds and covers a distance of 120 meters. What is the approximate acceleration of the car?

OpenStudy (anonymous):

v = u+at :initial velocity(u) = zero hence, v = at :eq1 we also know that: v^2 = u^2+2as :initial velocity(u) = zero v^2 = 2as :eq2 (at)^2 = 2as :v = at eq1. a = 2s/t^2 = 6m/s^2

OpenStudy (anonymous):

A.5 meters/second2 B.6 meters/second2 C.9 meters/second2 D.12 meters/second2 E.37 meters/second2